# Quantifying Azure RBAC Wildcard Overreach

### v2

@@ Abstract
Azure’s Role-Based Access Control (RBAC) leverages wildcard permissions to simplify policy authoring, but this abstraction often obscures the actual set of allowed operations and undermines least-privilege guarantees. We introduce Belshazaar, a two-stage framework that targets both (1) the effective permission set problem—deriving the exact list of actions granted by arbitrary wildcard specifications—and (2) the evaluation of wildcards permissions ”spread”. First, we formalize Azure action syntax via a context-free grammar and implement a compiler that expands any wildcard into its explicit action set. Second, we define an ultrametric diameter metric to quantify semantic overreach in wildcard scenarios. Applied to Microsoft’s official catalog of 15,481 actions, Belshazaar reveals that about 35% of actions admit a cross Resource Provider reach when associated with non obvious wildcards, and that effective permissions sets are effectively computable. These findings demonstrate that wildcard patterns can introduce substantial privilege bloat, and that our approach offers a scalable, semantics-driven path toward tighter, least-privilege RBAC policies in Azure environments.
@@ 1.1 Findings
  • A review of all Azure Actions also shows that, when combined with ill-placed wildcards, 35% have a potential to get an extreme (cross Resource Provider) reach.
@@ 6.3.1 Initial Population
For each official Azure action string, we first constructed an initial population of 50 candidate wildcard patterns. Wildcards were inserted at random positions within the action string, subject to syntactic constraints designed to avoid trivial global patterns such as * or Microsoft.*.
@@ 6.3.2 Genetic Algorithm Optimization
After generating the initial population, we applied a Genetic Algorithm (GA) to evolve these wildcard populations over 50 generations. Each population consisted of 50 individuals associated with one original Azure action string.
@@
    Mutation and Reproduction: New candidates were produced by randomly mutating either the start or end position of the wildcard insertion interval. Mutations involved shifting positions by a small random positive or negative offset. Each mutated candidate was validated against the original syntactic constraints before inclusion in the population.
This iterative process allowed the algorithm to explore the search space of wildcard placements while avoiding trivial or degenerate cases.
@@
The full evolutionary search was executed across all official Azure actions, requiring approximately 2 hours of computation on a standard desktop machine. The output consists of, for each Azure action, one or more wildcard patterns whose expansions exhibit the smallest ultrametric diameters. These extreme wildcard patterns represent dangerous generalizations in terms of permission breadth.
@@ 6.3.3 Distribution of Diameters
Because ultrametric diameters take on discrete integer values, we report the median via linear interpolation on the empirical cumulative distribution. After sorting the diameter counts and computing the cumulative percentages, we locate the two adjacent diameter values that straddle the 50% threshold and interpolate between them. This yields a more precise median of 1.32 rather than an integer.
@@ 6.3.4 Wide-ranging wildcards
Microsoft.Compu*egister/action Microsoft.Compute/register/action Microsoft.ComputeSchedule/register/action
@@
Microsoft.Con*ts/delete Microsoft.Confluent/organizations/environments/delete Microsoft.ConnectedOpenStack/networkPorts/delete
@@
Microsoft.Edg*/write Microsoft.Edge/capabilityLists/write Microsoft.EdgeMarketplace/locations/operationStatuses/write
@@
Table 3: A sample of diameter 2 pairs produced by our generator
@@ 8 Conclusion
In this paper, we presented Belshazaar, a semantics–aware framework for rigorously expanding and evaluating Azure RBAC wildcards. By formalizing the complete Azure action language as a context-free grammar and building a compiler to derive minimal explicit permission sets, Belshazaar delivers an accurate “effective permissions” view for any wildcard specification. Our linear ultrametric diameter model quantifies the semantic overreach of non-obvious wildcards, showing that approximately 35% span more than one Resource Provider.

### v3

@@ Abstract
Azure’s Role-Based Access Control (RBAC) leverages wildcard permissions to simplify policy authoring, but this abstraction often obscures the actual set of allowed operations and undermines least-privilege guarantees. We introduce Belshazaar, a two-stage framework that targets both (1) the effective permission set problem—deriving the exact list of actions granted by arbitrary wildcard specifications—and (2) the evaluation of wildcards permissions ”spread”. First, we formalize Azure action syntax via a context-free grammar and implement a compiler that expands any wildcard into its explicit action set. Second, we define an ultrametric diameter metric to quantify semantic overreach in wildcard scenarios. Applied to Microsoft’s official catalog of 15,481 actions, Belshazaar reveals that about 50% of actions admit a cross Resource Provider reach when associated with non obvious wildcards, and that effective permissions sets are effectively computable. These findings demonstrate that wildcard patterns can introduce substantial privilege bloat, and that our approach offers a scalable, semantics-driven path toward tighter, least-privilege RBAC policies in Azure environments.
@@ 1.1 Findings
  • A review of all Azure Actions also shows that, when combined with ill-placed wildcards, 50% have a potential to get an extreme (cross Resource Provider) reach.
@@ 6.3.1 Initial Population
For each official Azure action string, we first constructed an initial population of 40 candidate wildcard patterns. Wildcards were inserted at random positions within the action string, subject to syntactic constraints designed to avoid trivial global patterns such as * or Microsoft.*.
@@ 6.3.2 Genetic Algorithm Optimization
After generating the initial population, we applied a Genetic Algorithm (GA) to evolve these wildcard populations over 10 generations. Each population consisted of 40 individuals associated with one original Azure action string.
@@
    Mutation and Reproduction: New candidates were produced by randomly mutating either the start or end position of the wildcard insertion interval. Mutations involved shifting positions by a random positive or negative offset. Each mutated candidate was validated against the original syntactic constraints before inclusion in the population.
This iterative process allowed the algorithm to explore the search space of wildcard placements while avoiding trivial or degenerate cases, and to give a lower bound estimate of the number of extreme pairs.
@@
The full evolutionary search was executed across all official Azure actions, requiring approximately 3.5 hours of computation on a standard desktop machine. The output consists of, for each Azure action, one or more wildcard patterns whose expansions exhibit the smallest ultrametric diameters. These extreme wildcard patterns represent dangerous generalizations in terms of permission breadth.
@@ 6.3.3 Distribution of Diameters
Because ultrametric diameters take on discrete integer values, we report the median via linear interpolation on the empirical cumulative distribution. After sorting the diameter counts and computing the cumulative percentages, we locate the two adjacent diameter values that straddle the 50% threshold and interpolate between them. This yields a more precise median of 0.98 rather than an integer.
@@
Distribution of action pairs by diameter
Almost all Azure actions can be paired by non-obvious wildcards: about half of all pairs are cross Resource Provider (extreme pairs of diameter 1), another half are same provider, but without least privilege (diameter 2). The number of least privileged pairs (of diameter 3 and above) is less than 3%. No pairs of diameter greater than 5 were found.
Remember that the number of extreme pairs is a lower bound estimate.
@@ 6.3.4 Wide-ranging wildcards
Table 3: A sample of non-extreme (diameter 2) pairs produced by our generator
@@ 8 Conclusion
In this paper, we presented Belshazaar, a semantics–aware framework for rigorously expanding and evaluating Azure RBAC wildcards. By formalizing the complete Azure action language as a context-free grammar and building a compiler to derive minimal explicit permission sets, Belshazaar delivers an accurate “effective permissions” view for any wildcard specification. Our linear ultrametric diameter model quantifies the semantic overreach of non-obvious wildcards, showing that approximately 50% span more than one Resource Provider.